Calculate the number of vertices in a full 5-ary tree with 45 internal vertices.Also find out the number of leaves.


1
Expert's answer
2021-06-17T15:10:48-0400

A full "m" ‐ary tree with "i" internal vertices has "n=m\cdot i+1" vertices. 


"n=m\cdot i+1=5\cdot 45+1=226"

A full "m" ‐ary tree with "i" internal vertices has "l=(m-1)i+1" leaves.


"l=(m-1)i+1=(5-1)\cdot45+1=181"